OpenStudy (anonymous):

Find a • b. a = 4i + 8j, b = -2i + 3j <2, 11> 32 <-8, 24> 16

jimthompson5910 (jim_thompson5910):

Hint: If v = a*i + b*j and w = c*i + d*j then v dot w = a*c + b*d

OpenStudy (anonymous):

ok onesec let me calculate

OpenStudy (anonymous):

so its 16

jimthompson5910 (jim_thompson5910):

16 is correct